# Rectangular recurrence relations in $\mathfrak{gl}_{n}$ and $\mathfrak{o}_{2n+1}$ invariant integrable models

### v1

@@ 5 Conclusion
Using the zero modes method we describe in this paper the rectangular recurrence relations for the off-shell Bethe vectors in $\mathfrak{gl}_{n}$- and $\mathfrak{o}_{2n+1}$-invariant integrable models. These relations are presented as sums of the actions of the monodromy entries $T_{i,j}(z)$ with $i<j$ on the off-shell Bethe vectors ${\mathbb{B}}({\bar{t}})$ and sums over partitions of the sets of Bethe parameters. When one of the set ${\bar{t}}^{s}$ becomes empty the recurrence relations given by theorem 4.1 describe a factorization of the Bethe vectors and relates to the symmetry of $\mathfrak{gl}_{n}$-type Bethe vectors found in the paper [21]. Whether there exists a similar symmetry for $\mathfrak{o}_{2n+1}$-type Bethe vectors remains an open question. Such a symmetry would allow to relate the recurrence relation (4.24) to the recurrence relation (4.22). If it exists, this symmetry would require to explore the explicit relations between the monodromy matrix entries $T_{i,j}(z)$ following from the symmetry relation (2.9). It makes the search for a symmetry relation of $\mathfrak{o}_{2n+1}$-type monodromy matrix much more involved than in the $\mathfrak{gl}_{n}$ case.
@@ Acknowledgement
$\begin{split}\mathcal{Z}_{\ell}^{n}\cdot{\mathbb{B}}({\bar{t}})&=\frac{1}{\mu_{\ell}^{n}(z;{\bar{t}})}\ \sum_{i=1}^{\ell}\sum_{\rm part}g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\ \prod_{p=i}^{\ell-1}\Omega^{R}({\bar{t}}^{p}_{\scriptscriptstyle\rm I},{\bar{t}}^{p}_{\scriptscriptstyle\rm II}|{\bar{t}}^{p-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{p+1}_{\scriptscriptstyle\rm II})\times\\ &\quad\times T_{i,n}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell-1},\{{\bar{t}}^{s}\}_{\ell}^{n-1})\,,\end{split}$ (A.1)
where the sum goes over partitions ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}$ with cardinalities $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=1$ for all $s=i,\ldots,\ell-1$ while ${\bar{t}}^{\ell}$ is not parted with ${\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\equiv{\bar{t}}^{\ell}$.
@@
Using the induction assumption, relation (A.2) rewrites
@@
$\begin{split}-\chi_{\ell}\ T_{\ell+1,n}(z)\cdot{\mathbb{B}}({\bar{t}})&+\sum_{i=1}^{\ell}\sum_{\rm part}g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\prod_{p=i}^{\ell-1}\Omega^{R}({\bar{t}}^{p}_{\scriptscriptstyle\rm I},{\bar{t}}^{p}_{\scriptscriptstyle\rm II}|{\bar{t}}^{p-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{p+1}_{\scriptscriptstyle\rm II})\Big{|}_{{\bar{t}}^{\ell}\vdash\{{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\}}\times\\ &\times\Big{(}\chi_{\ell+1}\,\alpha_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\,\Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})-\chi_{\ell}\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})\Big{)}\times\\ &\qquad\times T_{i,n}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell},\{{\bar{t}}^{s}\}_{\ell+1}^{n-1})\Big{)}\,.\end{split}$ (A.5)
@@
$\begin{split}&\Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|\{{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II}\},{\bar{t}}^{\ell+1})=h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\ \Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})\\ &\Omega^{R}({\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-2}_{\scriptscriptstyle\rm II},\{{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\})=h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})\end{split}$
@@
$\Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|\{{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II}\},{\bar{t}}^{\ell+1})=\frac{1}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}\ \Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})$
we conclude that the equality between (A.3) and (A.5) is equivalent to the relation (A.1) for $\ell\to\ell+1$ due to the identity
$\begin{split}&g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\left(h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})-\frac{f({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}\right)=\frac{g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}\Big{(}f({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})-f({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)\Big{)}=\\ &\quad=\frac{g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}\Big{(}g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})-g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)\Big{)}=\frac{g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\,g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\,g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\,g({\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},z)}=g(z,{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\,.\end{split}$ (A.6)
This ends the inductive proof of relation (A.1). $\square$
@@
$\begin{split}&\mu^{k+1}_{\ell}(z;{\bar{t}})\ {\sf T}_{k+1,k}\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},\{\bar{u}^{s}\}_{\ell}^{k},\{{\bar{t}}^{s}\}_{k+1}^{n-1})=\chi_{k+1}\,\mu^{k}_{\ell}(z;{\bar{t}})\,\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})\ +\\ &\quad+\sum_{i=1}^{\ell}\sum_{j=k+1}^{n}\sum_{\rm part}\Big{(}\chi_{k+1}\,\alpha_{k}({\bar{t}}^{k}_{\scriptscriptstyle\rm III})\,\Omega^{L}({\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}|{\bar{t}}^{k-1},{\bar{t}}^{k+1})\,f(z,{\bar{t}}^{k}_{\scriptscriptstyle\rm III})-\chi_{k}\,\Omega^{R}({\bar{t}}^{k}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm II}|{\bar{t}}^{k-1},{\bar{t}}^{k+1})\Big{)}\times\\ &\quad\times\Xi^{\ell,k+1}_{i,j}(z;{\bar{t}})\ T_{i,j}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell-1},\{{\bar{t}}^{s}\}_{\ell}^{k-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{k}^{j-1},\{{\bar{t}}^{s}\}_{j}^{n-1})\,.\end{split}$ (A.7)
@@
$\begin{split}\Xi^{\ell,k+1}_{i,k+1}(z;{\bar{t}})=\Xi^{\ell,k}_{i,k}(z;{\bar{t}})&=\psi_{\ell}(z;{\bar{t}}_{\scriptscriptstyle\rm I})\prod_{p=i}^{\ell-1}\Omega^{R}({\bar{t}}^{p}_{\scriptscriptstyle\rm I},{\bar{t}}^{p}_{\scriptscriptstyle\rm II}|{\bar{t}}^{p-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{p+1}_{\scriptscriptstyle\rm II})\,,\\ \Xi^{\ell,k+1}_{i,j}(z;{\bar{t}})\Big{|}_{{\bar{t}}^{k}\vdash\{{\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}\}}&=h({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})\ \Xi^{\ell,k+1}_{i,j}(z;{\bar{t}})\Big{|}_{{\bar{t}}^{k}\vdash\{{\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}\}}\,,\\ \Omega^{R}({\bar{t}}^{k}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm II}|{\bar{t}}^{k-1},{\bar{t}}^{k+1})\Big{|}_{{\bar{t}}^{k+1}\vdash\{{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm III}\}}&=h({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})\ \Omega^{R}({\bar{t}}^{k}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm II}|{\bar{t}}^{k-1},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II})\,,\\ \Omega^{L}({\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}|{\bar{t}}^{k-1},{\bar{t}}^{k+1})\Big{|}_{{\bar{t}}^{k+1}\vdash\{{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm III}\}}&=\frac{1}{g({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})}\ \Omega^{L}({\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}|{\bar{t}}^{k-1},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II})\end{split}$
@@
$\phi_{k+1}(z;{\bar{t}}_{\scriptscriptstyle\rm III})\Big{(}h({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})-\frac{f(z,{\bar{t}}^{k}_{\scriptscriptstyle\rm iii})}{g({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})}\Big{)}=\phi_{k}(z,{\bar{t}}_{\scriptscriptstyle\rm III})$
the equality between (A.7) and (A.8) is equivalent to (3.10). This finishes the inductive proof of the proposition 3.1. $\square$
@@
where the sum goes over partitions ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III}\}$ with cardinalities $|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=1$ for all $s=k,\ldots,j-1$ while the set ${\bar{t}}^{k-1}$ is not parted and ${\bar{t}}^{k-1}_{\scriptscriptstyle\rm II}\equiv{\bar{t}}^{k-1}$.

### v2

@@ 5 Conclusion
Using the zero modes method we describe in this paper the rectangular recurrence relations for the off-shell Bethe vectors in $\mathfrak{gl}_{n}$- and $\mathfrak{o}_{2n+1}$-invariant integrable models. These relations are presented as sums of the actions of the monodromy entries $T_{i,j}(z)$ with $i<j$ on the off-shell Bethe vectors ${\mathbb{B}}({\bar{t}})$ and sums over partitions of the sets of Bethe parameters. When one of the set ${\bar{t}}^{s}$ becomes empty the recurrence relations given by theorem 4.1 describe a factorization of the Bethe vectors and relates to the symmetry of $\mathfrak{gl}_{n}$-type Bethe vectors found in the paper [22].
@@ Acknowledgement
$\begin{split}\mathcal{Z}_{\ell}^{n}\cdot{\mathbb{B}}({\bar{t}})=\frac{1}{\mu_{\ell}^{n}(z;{\bar{t}})}\ \sum_{i=1}^{\ell}\sum_{\rm part}&\ g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\ \prod_{p=i}^{\ell-1}\Omega^{R}({\bar{t}}^{p}_{\scriptscriptstyle\rm I},{\bar{t}}^{p}_{\scriptscriptstyle\rm II}|{\bar{t}}^{p-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{p+1}_{\scriptscriptstyle\rm II})\times\\ &\times T_{i,n}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell-1},\{{\bar{t}}^{s}\}_{\ell}^{n-1})\,,\end{split}$ (A.1)
where the sum goes over partitions ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}$ with cardinalities $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=1$ for all $s=i,\ldots,\ell-1$, ${\bar{t}}^{\ell}$ is not parted with ${\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\equiv{\bar{t}}^{\ell}$, ${\bar{t}}^{i-1}_{\scriptscriptstyle\rm II}={\bar{t}}^{i-1}$, ${\bar{t}}^{i-1}_{\scriptscriptstyle\rm I}=\varnothing$, and ${\bar{t}}^{0}_{\scriptscriptstyle\rm I}=\varnothing$.
@@
Here the sum goes over partitions ${\bar{t}}^{\ell}\vdash\{{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\}$ with cardinality $|{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|=1$. In the last line of (A.2) we wrote explicitly the arguments of the function $\mu^{n}_{\ell}(z;{\bar{t}}^{\ell-1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{n-1})$ for clarification. Using the induction assumption, relation (A.2) rewrites
@@
Here the sum goes over partitions ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}$ with cardinalities $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=1$ for all $s=i,\ldots,\ell$, ${\bar{t}}^{i-1}_{\scriptscriptstyle\rm II}={\bar{t}}^{i-1}$, and ${\bar{t}}^{i-1}_{\scriptscriptstyle\rm I}=\varnothing$.
@@
$\begin{split}-\chi_{\ell}\ T_{\ell+1,n}(z)\cdot{\mathbb{B}}({\bar{t}})&+\sum_{i=1}^{\ell}\sum_{\rm part}g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\prod_{p=i}^{\ell-2}\Omega^{R}({\bar{t}}^{p}_{\scriptscriptstyle\rm I},{\bar{t}}^{p}_{\scriptscriptstyle\rm II}|{\bar{t}}^{p-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{p+1}_{\scriptscriptstyle\rm II})\ \Omega^{R}({\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-2}_{\scriptscriptstyle\rm II},\{{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\})\times\\ &\times\Big{(}\chi_{\ell+1}\,\alpha_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\,\Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})-\chi_{\ell}\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})\Big{)}\times\\ &\qquad\times T_{i,n}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell},\{{\bar{t}}^{s}\}_{\ell+1}^{n-1})\Big{)}\,.\end{split}$ (A.5)
Here the sum over partitions is the same as in (A.3).
@@
$\begin{split}&\Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|\{{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II}\},{\bar{t}}^{\ell+1})=h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\ \Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})\\ &\Omega^{R}({\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-2}_{\scriptscriptstyle\rm II},\{{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\})=h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\ \Omega^{R}({\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-2}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II})\end{split}$
@@
$\Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|\{{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II}\},{\bar{t}}^{\ell+1})=\frac{1}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})$
we conclude that the equality between (A.3) and (A.5) is equivalent to the relation (A.1) at $\ell\to\ell+1$ due to the identities
$\begin{split}&g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\left(h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})-\frac{f({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}\right)=\frac{g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}\Big{(}f({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})-f({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)\Big{)}=\\ &\quad=\frac{g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}\Big{(}g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})-g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)\Big{)}=\frac{g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\,g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\,g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\,g({\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},z)}=g(z,{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\,\end{split}$ (A.6)
for $i<\ell$ and $1-f({\bar{t}}^{i}_{\scriptscriptstyle\rm I},z)=g(z,{\bar{t}}^{i}_{\scriptscriptstyle\rm I})$ for $i=\ell$. This ends the inductive proof of relation (A.1). $\square$
@@
$\begin{split}&\mu^{k+1}_{\ell}(z;{\bar{t}})\ {\sf T}_{k+1,k}\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},\{\bar{u}^{s}\}_{\ell}^{k},\{{\bar{t}}^{s}\}_{k+1}^{n-1})=\chi_{k+1}\,\mu^{k}_{\ell}(z;{\bar{t}})\,\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})\ +\\ &\quad+\sum_{i=1}^{\ell}\sum_{j=k+1}^{n}\sum_{\rm part}\Big{(}\chi_{k+1}\,\alpha_{k}({\bar{t}}^{k}_{\scriptscriptstyle\rm III})\,\Omega^{L}({\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}|{\bar{t}}^{k-1},{\bar{t}}^{k+1})\,f(z,{\bar{t}}^{k}_{\scriptscriptstyle\rm III})-\chi_{k}\,\Omega^{R}({\bar{t}}^{k}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm II}|{\bar{t}}^{k-1},{\bar{t}}^{k+1})\Big{)}\times\\ &\quad\times\Xi^{\ell,k+1}_{i,j}(z;\{{\bar{t}}^{s}\}_{1}^{k-1},{\bar{t}}^{k}_{\scriptscriptstyle\rm II},\{{\bar{t}}^{s}\}_{k+1}^{n-1})\ T_{i,j}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell-1},\{{\bar{t}}^{s}\}_{\ell}^{k-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{k}^{j-1},\{{\bar{t}}^{s}\}_{j}^{n-1})\,.\end{split}$ (A.7)
@@
$\begin{split}\Xi^{\ell,k+1}_{i,k+1}(z;{\bar{t}})=\Xi^{\ell,k}_{i,k}(z;{\bar{t}})&=g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\prod_{p=i}^{\ell-1}\Omega^{R}({\bar{t}}^{p}_{\scriptscriptstyle\rm I},{\bar{t}}^{p}_{\scriptscriptstyle\rm II}|{\bar{t}}^{p-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{p+1}_{\scriptscriptstyle\rm II})\,,\\ \Xi^{\ell,k+1}_{i,j}(z;{\bar{t}})\Big{|}_{{\bar{t}}^{k}\vdash\{{\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}\}}&=h({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})\ \Xi^{\ell,k+1}_{i,j}(z;\{{\bar{t}}^{s}\}_{1}^{k-1},{\bar{t}}^{k}_{\scriptscriptstyle\rm II},\{{\bar{t}}^{s}\}_{k+1}^{n-1}),\\ \Omega^{R}({\bar{t}}^{k}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm II}|{\bar{t}}^{k-1},\{{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm III}\})&=h({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})\ \Omega^{R}({\bar{t}}^{k}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm II}|{\bar{t}}^{k-1},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II})\,,\\ \Omega^{L}({\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}|{\bar{t}}^{k-1},\{{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm III}\})&=\frac{1}{g({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})}\ \Omega^{L}({\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}|{\bar{t}}^{k-1},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II})\end{split}$
@@
$g({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},z)\Big{(}h({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})-\frac{f(z,{\bar{t}}^{k}_{\scriptscriptstyle\rm III})}{g({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})}\Big{)}=g({\bar{t}}^{k}_{\scriptscriptstyle\rm III},z)$
in the equality between (A.7) and (A.8) coefficients at $\chi_{k}$ cancel each other and coefficients at $\chi_{k+1}$ yield (3.10). This finishes the inductive proof of the proposition 3.1. $\square$
@@
where the sum goes over partitions ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III}\}$ with cardinalities $|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=1$ for all $s=k,\ldots,j-1$, the set ${\bar{t}}^{k-1}$ is not parted, ${\bar{t}}^{k-1}_{\scriptscriptstyle\rm II}\equiv{\bar{t}}^{k-1}$, and ${\bar{t}}^{n}_{\scriptscriptstyle\rm III}=\varnothing$.